OpenStudy (anonymous):

A 95-kilogram boy is running at a speed of 7 m/s. What is the boy’s kinetic energy? *

OpenStudy (cggurumanjunath):

K.E. = 0.5 * m * v ^2

OpenStudy (cggurumanjunath):

k.e.= 0.5*95*7*7 =2327.5J OR 2.375 KJ